simplify and write answers without exponents
125^-2/3
(1/8)^2/3

thank you so much for your time and effort!

Answer by rothauserc(4718)

A) 125^(-2/3) = 1 / 125^(2/3) = 1/ 15625^(1/3) = 1 / 25 = 0.04
B) (1/8)^2/3 = (1 / 64)^(1/3) = 1 / 4 = 0.25